Static Equilibrium of a ParticleA particle is subjected to three forces, F_1, F_2
Answer #1

F1 = 9 N
F2 = 10 N
α = 57
β = 59

To Maintain Static Equilibrium,
F1 * cos(α) + F3 * cos(γ) = F2 * sin(β)
9 * cos(57) + F3 * cos(γ) = 10 * sin(59)
F3 * cos(γ) = 3.67 -------------1

F1 * sin(α) + F3 * sin(γ) = F2 * cos(β)
9 * sin(57) + F3 * sin(γ) = 10 * cos(59)
F3 * sin(γ) = - 2.397 -------------2

Dividing eq 2/1
tan(γ) = - 2.397/3.67
γ = tan^-1(- 2.397/3.67)
γ = - 33.15o

F3 * cos(-33.15) = 3.67
F3 = 4.4 N